# Detectability of Lensed Gravitational Waves in Matched-Filtering Searches

### v1

@@ Abstract
Gravitational lensing by compact, small-scale intervening masses causes frequency-dependent distortions to gravitational-wave events. The optimal signal-to-noise ratio (SNR) is often used as a proxy for the detectability of exotic signals in gravitational-wave searches. In reality, the detectability of such signals in a matched-filtering search requires comprehensive consideration of match-filtered SNR, signal-consistency test value, and other factors. In this work, we investigate for the first time the detectability of lensed gravitational waves from compact binary coalescences with a match-filtering search pipeline, GstLAL. Contrary to expectations from the optimal-SNR approximation approach, we show that the strength of a signal (i.e., higher optimal SNR) does not necessarily result in higher detectability. We also demonstrate that lensed gravitational waves with wave optics effects can suffer significantly, from $\leavevmode\nobreak\ 90\%$ (unlensed) to $<1\%$ (lensed) detection efficiency, due to downranking by the signal-consistency test values. These findings stress the need to extend current template banks to effectively search for lensed gravitational waves and to reassess current constraints on compact dark matter scenarios.
@@ I Introduction
Gravitational wave detection typically relies on searching for signals using templates based on General Relativity (GR) Allen et al. (2012). Commonly, templates used in the modelled search consider a simple quasi-circular quadrupole mode waveform Usman et al. (2016); Abbott et al. (2019); Nitz et al. (2019); Venumadhav et al. (2019, 2020); Abbott et al. (2021a); Nitz et al. (2020); Abbott et al. (2024, 2023a); Andres et al. (2022); Nitz et al. (2021, 2023); Olsen et al. (2022); Mehta et al. (2023); Kumar and Dent (2024). While spin-orbit precessing effect, higher-order mode, and tidal deformability have been considered recenrly Wadekar et al. (2023); Schmidt et al. (2024a, b); Schmidt and Caudill (2024); Wadekar et al. (2024); Chia et al. (2024), there can be other physical effects such as orbital eccentricity, enviromental effects and deviations from GR. Within the GR framework, gravitational lensing emerges as the dominant mechanism for modifying gravitational waves during their propagation through space Deguchi and Watson (1986); Nakamura (1998); Takahashi and Nakamura (2003). A distorted gravitational wave signal, either by lensing or other means, can be problematic to detect with incomplete templates in modelled searches, potentially leading to missed detections or mischaracterizations of the source properties. To determine whether the signals are detectable, one must know the probability that a random noise fluctuation produces the same effects in modelled searches, i.e., the false-alarm-rate, which typical optimal sigal-to-noise ratio (SNR) analyses cannot capture.
@@
Gravitational waves can be deflected and distorted by intervening massive objects. Consider the primary gravitational wave sources for the ground-based detector, characterized by wavelength $\lambda_{\rm{GW}}\sim 10^{6}$ m, and a configuration in which the observer and the source are well aligned with the lens. Then, if the characteristic size of the lens $R_{\rm{Lens}}\gg\lambda_{\rm{GW}}$, such as galaxies ($R_{\rm{Lens}}\sim 10^{20}$ m) or galaxy clusters ($R_{\rm{Lens}}\sim 10^{23}$ m), gravitational waves follow well-defined paths, and their magnification is achromatic Wang et al. (1996); Dai and Venumadhav (2017); Ezquiaga et al. (2021). This leads to distinct and well-separated (in time) copies of the original chirp that differ by a constant phase shift, relative magnification, and a relative time delay. This is the regime known as strong lensing, where multiple images are produced. When $R_{\rm{Lens}}\sim\lambda_{\rm{GW}}$, lenses such as intermediate-mass black holes and dark matter clumps within a galaxy can cause a frequency-dependent modulation of the waveform. This modulation, involving both interference and diffraction effects, can leave a lensing-induced distortion on a single signal Deguchi and Watson (1986); Nakamura (1998); Takahashi and Nakamura (2003); Leung et al. (2023). These distortions hold invaluable information about the possible population of small compact lenses Jung and Shin (2019); Lai et al. (2018); Christian et al. (2018); Dai et al. (2018); Diego et al. (2019); Diego (2020); Liao et al. (2020); Mishra et al. (2021); Liu et al. (2023); Chakraborty and Mukherjee (2024), the caustic structure of different lenses Lo et al. (2024), the halo mass function Fairbairn et al. (2023); Tambalo et al. (2023); Çalışkan et al. (2023); Savastano et al. (2023), and the cosmic expansion rate Cremonese et al. (2021); Chen et al. (2024). Understanding them is also key to avoid biases in parameter estimation Cao et al. (2014) and false claims of violations of general relativity Ezquiaga et al. (2020); Gupta et al. (2024).
@@
Since the LIGO-Virgo-KAGRA (LVK) collaboration Aasi et al. (2015); Acernese et al. (2015) first successful detection of gravitational waves in 2015, efforts have been continuously made to search for lensing signatures within gravitational-wave data Hannuksela et al. (2019); Abbott et al. (2021b, 2023b); Janquart et al. (2023, 2024). No significant evidence has been found yet. Other efforts pointed that no lensing features have been detected McIsaac et al. (2020); Li et al. (2023); Liu et al. (2021), although Dai et al. Dai et al. (2020) also highlighted an intriguing candidate that is however astrophysically unlikely. These analyses are performed after detecting gravitational waves by different search algorithms designed to search for (unlensed) gravitational waves from compact binary coalescences (CBCs) Allen et al. (2012); Abbott et al. (2020). The search algorithms use a pre-defined template bank to detect all the signals from CBCs Abbott et al. (2019, 2021a, 2023a). However, as we will show in this work, signals can be missed if the lensing distortion is large. This can lead to biases in follow-up lensing analyses. The probability of detecting a lensed signal thus may depend on the source and lens parameters, making it crucial to examine. As the current templates lack gravitational lensing physics, it is fundamental to understand its impact on detectability in matched-filtering searches.
@@
In this work, we study for the first time the detectability of lensed signals with state-of-the-art gravitational-wave search pipelines. We perform injection campaigns [^1] In gravitational-wave searches, to test the search efficiency of a pipeline towards a certain type of gravitational-wave signals, we inject a set of simulated signals in real noise and assess how many signals the pipeline can recover. This is known as an injection campaign. to systematically study how lensing influences the detectability of gravitational-wave signals using a time-domain matched-filtering search pipeline GstLAL Messick et al. (2017); Sachdev et al. (2019); Cannon et al. (2021); Tsukada et al. (2023). GstLAL is one of the main pipelines the LVK collaboration uses to search for gravitational-waves signals. We examine the accuracy of the recovered source parameters, matched-filter SNRs, the signal consistency test values, and the false-alarm rates of the injections.
@@
This paper is structured as follows: Section II introduces the fundamentals of matched-filtering. Section III describes the essential parts of matched-filtering in GstLAL. Section IV introduces gravitational lensing of gravitational waves using the point mass lens model and illustrates the expected effects of lensing on the detectability. Section V describes the injection sets and the setup of the matched-filtering searches in our study. Section VI shows the results of the role of lensing and background noises in the detectability of lensed gravitational waves. Section VII concludes the results and proposes new strategies for detecting lensed signatures in gravitational waves and underscores the implication of our work.
II Matched-Filtering Search for Gravitational wave
@@
To access the detectability of gravitational-waves signals, one needs to predict their SNR in matched-filtering searches. Here we introduce the fundamentals of matched-filtering search and discuss the current efforts of detectability studies.
@@ II.1 Overview of Matched-filtering Search
The expected gravitational-wave signals from CBC are well characterized by combinations of the binary component parameters. Extrinsic parameters, such as source distances and inclinations, are not included in the template of matched-filtering searches. In the case of the matched-filtering search conducted in this study, which considers the intrinsic parameters, the parameter space is defined by the binary’s component masses in the detector frame and spins. The component mass is denoted as $m_{i}$, where $i=\{1,2\}$, and by convention, $m_{1}\geq m_{2}$. The dimensionless spin is given as $\vec{\chi}_{i}=c\vec{S}_{i}/Gm^{2}_{i}$, where $c$ is the speed of light, and $G$ is the gravitational constant, and $\vec{S}_{i}$ is the spin angular momentum.
@@
Matched-filtering is a powerful technique that is used to search for gravitational waves from compact binary coalescences and it is the optimal filter for detecting waveforms in stationary Gaussian noise Allen et al. (2012). For a given search parameter space (e.g. in the component masses $m_{1}$-$m_{2}$ space), a set of gravitational waveforms $h_{T,\;j}$ (also known as “templates”) is precomputed and collectively stored in a “template bank”. The parameters of the templates are $\vec{\theta}=\{m1,m2,\chi_{1},\chi_{2}\}$ and in the detector frame Krolak and Schutz (1987).
To identify a real gravitational-wave signal $h$, we compute the correlation between the template waveform $h_{T,\;j}$ and the data $d$ (which is composed of the background noise and potentially a signal) to construct the signal-to-noise-ratio (SNR) $\rho$. In the time domain, this is  Allen et al. (2012)
@@
The SNR $\rho(t)$ is defined as the modulus of the complex SNR, i.e. $\rho(t)=|z_{j}(t)|$.
@@
Figure 1: Left: Match between lensed signals and unlensed templates as a function of the (redshifted) lens mass $M_{\rm Lz}$ and impact parameter $y$. Right: Ratio of the optimal signal-to-noise ratio of lensed ($\rho_{\rm L}$ and unlensed ($\rho_{\rm UL}$) signals for the same $(M_{\rm Lz},y)$ parameter space). Here, we use a spinless binary black hole signal at 125 Mpc with $m_{1}=m_{2}=30M_{\odot}$. $\rho_{\rm{L}}/\rho_{\rm{UL}}>1$ holds true for point mass lens model.
@@ II.2 Current Efforts in Detectability Study
In addition, the phase and time shifts, $\phi_{0}$ and $\tau_{0}$, maximize the match between the template waveform $h_{T,\;j}$ and the signal $h$. High $\mathcal{M}$ values indicate that $h_{T,\;j}$ and $h$ are well-matched, where a maximum value of $\mathcal{M}=1$ indicates that the signal is exactly described by the template waveform. In the case where the template waveform exactly matches the signal (i.e. $h=h_{T,\;j}$) and the SNR is optimized with respect to noise realizations, the SNR is called the “optimal SNR” $\rho_{\rm opt}$ , and the “observed SNR” $\rho_{\rm obs}$ otherwise.
Various methods, each with its own strengths and limitations, have been used to assess gravitational-wave signal detectability. One widely used approach, employed in studies by Tiwari (2018); Veske et al. (2021); Gerosa et al. (2020), calculates the optimal signal-to-noise ratio (SNR) from a network of detectors, assuming that the same source parameters are recovered in matched-filtering searches. This method has been applied to evaluate the detectability of gravitational-wave signals with wave optics effects, as seen in works by Bondarescu et al. (2023); Mishra et al. (2024). While this assumption isn’t strictly valid due to noise fluctuations, it is a useful proxy for assessing detectability, particularly in high SNR scenarios.  Essick (2023) improved this method by accounting for stationary Gaussian noise fluctuations, providing a more realistic estimate of detectability. Although these approaches don’t incorporate signal consistency tests, they offer a computationally efficient solution for analyzing millions of signals.
Magee el al. Magee et al. (2024) included the signal consistency test to investigate detectability. While their study did not compute complete likelihood-ratio ranking statistics, it did consider the impact of signal consistency test values on matched-filtering searches, offering valuable insights into this aspect of signal detection.
For the most comprehensive analysis of detectability, full injection campaigns are necessary. These campaigns simulate the detection process using search pipelines and simulated gravitational-wave signals. This approach provides detailed information about the injected signals, including matched-filter SNR, signal consistency test values, and final detection statistics. While computationally intensive, injection campaigns offer the most accurate and complete assessment of gravitational-wave signal detectability across various scenarios and detector configurations Messick et al. (2017).
@@ 6 Unlensed 2,3 1250 6217
Table 1: Set of 11 injections used in our three simulation campaigns. Injection campaign 1 studies the impact of lensing on detectability. Injection campaign 2 studies the role of signal strength. Injection campaign 3 gives a general picture of the detectability of the lensed gravitational waves. The found injection is defined as the false-alarm rate (FAR) of the injection below the FAR threshold $3.85\times 10^{-7}$ Hz (See Sec. III.2 for FAR definition). All injections used the $30-30M_{\odot}$ spinless binary black hole waveform (See Sec. V.2 for details on other source parameters.).

### v2

@@ Abstract
Gravitational lensing by compact, small-scale intervening masses causes frequency-dependent distortions to gravitational-wave events. The optimal signal-to-noise ratio (SNR) is often used as a proxy for the detectability of exotic signals in gravitational-wave searches. In reality, the detectability of such signals in a matched-filtering search requires comprehensive consideration of match-filtered SNR, signal-consistency test value, and other factors. In this work, we investigate for the first time the detectability of lensed gravitational waves from compact binary coalescences with a match-filtering search pipeline, GstLAL. Contrary to expectations from the optimal SNR approximation approach, we show that the strength of a signal (i.e., higher optimal SNR) does not necessarily result in higher detectability. We also demonstrate that lensed gravitational waves with wave optic effects can suffer significantly, from $\leavevmode\nobreak\ 90\%$ (unlensed) to $<1\%$ (lensed) detection efficiency, due to downranking by the signal-consistency test values. These findings stress the need to extend current template banks to effectively search for lensed gravitational waves and to reassess current constraints on compact dark matter scenarios.
@@ I Introduction
Gravitational-wave detection typically relies on searching for signals using templates based on general relativity (GR) Allen et al. (2012). Commonly, templates used in the modeled search consider a simple quasicircular quadrupole mode waveform Usman et al. (2016); Abbott et al. (2019); Nitz et al. (2019); Venumadhav et al. (2019, 2020); Abbott et al. (2021a); Nitz et al. (2020); Abbott et al. (2024, 2023a); Andres et al. (2022); Nitz et al. (2021, 2023); Olsen et al. (2022); Mehta et al. (2023); Kumar and Dent (2024). While spin-orbit precessing effect, higher-order mode, and tidal deformability have been considered recently Wadekar et al. (2023); Schmidt et al. (2024a, b); Schmidt and Caudill (2024); Wadekar et al. (2024); Chia et al. (2024), there can be other physical effects such as orbital eccentricity, environmental effects and deviations from GR. Within the GR framework, gravitational lensing emerges as the dominant mechanism for modifying gravitational waves during their propagation through space Deguchi and Watson (1986); Nakamura (1998); Takahashi and Nakamura (2003). A distorted gravitational-wave signal, either by lensing or other means, can be problematic to detect with incomplete templates in modeled searches, potentially leading to missed detections or mischaracterizations of the source properties. Gravitational lensing affects specific sources based on their alignment with an intervening mass, whereas non-GR effects generically alter the waveform of all gravitational wave sources regardless of their location or propagation path. To determine whether the signals are detectable, one must know the probability that a random noise fluctuation produces the same effects in modeled searches, i.e., the false-alarm rate, which typical optimal SNR analyses cannot capture.
@@
Gravitational waves can be deflected and distorted by intervening massive objects. Consider the primary gravitational-wave(GW) sources for the ground-based detector, characterized by wavelength $\lambda_{\rm{GW}}\sim 10^{6}$ m, and a configuration in which the observer and the source are well aligned with the lens. Then, if the characteristic size of the lens $R_{\rm{Lens}}\gg\lambda_{\rm{GW}}$, such as galaxies ($R_{\rm{Lens}}\sim 10^{20}$ m) or galaxy clusters ($R_{\rm{Lens}}\sim 10^{23}$ m), gravitational waves follow well-defined paths, and their magnification is achromatic Wang et al. (1996); Dai and Venumadhav (2017); Ezquiaga et al. (2021). This leads to distinct and well-separated (in time) copies of the original chirp that differ by a constant phase shift, relative magnification, and a relative time delay. This is the regime known as strong lensing, where multiple images are produced. When $R_{\rm{Lens}}\sim\lambda_{\rm{GW}}$, lenses such as intermediate-mass black holes and dark matter clumps within a galaxy can cause a frequency-dependent modulation of the waveform. This modulation, involving both interference and diffraction effects, can leave a lensing-induced distortion on a single signal Deguchi and Watson (1986); Nakamura (1998); Takahashi and Nakamura (2003); Leung et al. (2023); Liu et al. (2024). These distortions hold invaluable information about the possible population of small compact lenses Jung and Shin (2019); Lai et al. (2018); Christian et al. (2018); Dai et al. (2018); Diego et al. (2019); Diego (2020); Liao et al. (2020); Mishra et al. (2021); Liu et al. (2023); Chakraborty and Mukherjee (2024), the caustic structure of different lenses Lo et al. (2024), the halo mass function Fairbairn et al. (2023); Tambalo et al. (2023); Çalışkan et al. (2023); Savastano et al. (2023), and the cosmic expansion rate Cremonese et al. (2021); Chen et al. (2024). Understanding them is also key to avoid biases in parameter estimation Cao et al. (2014); Liu et al. (2024) and false claims of violations of GR Ezquiaga et al. (2020); Gupta et al. (2024); Liu et al. (2024).
@@
Since the LIGO-Virgo-KAGRA (LVK) Collaboration’s Aasi et al. (2015); Acernese et al. (2015) first successful detection of gravitational waves in 2015, efforts have been continuously made to search for lensing signatures within gravitational-wave data Hannuksela et al. (2019); Abbott et al. (2021b, 2023b); Janquart et al. (2023, 2024). No significant evidence has been found yet. Other efforts pointed out that no lensing features have been detected McIsaac et al. (2020); Li et al. (2023); Liu et al. (2021), although Dai et al. Dai et al. (2020) also highlighted an intriguing candidate that is, however, astrophysically unlikely. These analyses are performed after detecting gravitational waves by different search algorithms designed to search for (unlensed) gravitational waves from compact binary coalescences (CBCs) Allen et al. (2012); Abbott et al. (2020). The search algorithms use a predefined template bank to detect all the signals from CBCs Abbott et al. (2019, 2021a, 2023a). However, as we will show in this work, signals can be missed if the lensing distortion is large. This can lead to biases in follow-up lensing analyses. The probability of detecting a lensed signal thus may depend on the source and lens parameters, making it crucial to examine. As the current templates lack gravitational lensing physics, it is fundamental to understand its impact on detectability in matched-filtering searches.
@@
In this work, we study for the first time the detectability of lensed signals with state-of-the-art gravitational-wave search pipelines. We perform injection campaigns [^1] In gravitational-wave searches, to test the search efficiency of a pipeline toward a certain type of gravitational-wave signals, we inject a set of simulated signals in real noise and assess how many signals the pipeline can recover. This is known as an injection campaign. to systematically study how lensing influences the detectability of gravitational-wave signals using a time-domain matched-filtering search pipeline GstLAL Messick et al. (2017); Sachdev et al. (2019); Cannon et al. (2021); Tsukada et al. (2023). GstLAL is one of the main pipelines the LVK Collaboration uses to search for gravitational-wave signals. We examine the accuracy of the recovered source parameters, matched-filter SNRs, the signal-consistency test values, and the false-alarm rates (FARs) of the injections.
@@
This paper is structured as follows: Section II introduces the fundamentals of matched-filtering. Section III describes the essential parts of matched filtering in GstLAL. Section IV introduces gravitational lensing of gravitational waves using the point mass lens model and illustrates the expected effects of lensing on the detectability. Section V describes the injection sets and the setup of the matched filtering searches in our study. Section VI shows the results of the role of lensing and background noises in the detectability of lensed gravitational waves. Section VII concludes the results and proposes new strategies for detecting lensed signatures in gravitational waves and underscores the implication of our work.
II Matched-Filtering Search for Gravitational waves
@@
To access the detectability of gravitational-wave signals, one needs to predict their SNR in matched-filtering searches. Here we introduce the fundamentals of matched-filtering search and discuss the current efforts of detectability studies.
@@ II.1 Overview of matched-filtering search
The expected gravitational-wave signals from CBC are well characterized by combinations of the binary component parameters. Extrinsic parameters, such as source distances and inclinations, are not included in the template of matched-filtering searches. In the case of the matched-filtering search conducted in this study, which considers the intrinsic parameters, the parameter space is defined by the binary’s component masses in the detector frame and spins. The component mass is denoted as $m_{i}$, where $i=\{1,2\}$, and by convention, $m_{1}\geq m_{2}$. The dimensionless spin is given as $\vec{\chi}_{i}=c\vec{S}_{i}/Gm^{2}_{i}$, where $c$ is the speed of light, $G$ is the gravitational constant, and $\vec{S}_{i}$ is the spin angular momentum.
@@
Matched filtering is a powerful technique that is used to search for gravitational waves from CBCs and it is the optimal filter for detecting waveforms in stationary Gaussian noise Allen et al. (2012). For a given search parameter space (e.g. in the component masses $m_{1}$-$m_{2}$ space), a set of gravitational waveforms $h_{T,\;j}$ (also known as “templates”) is precomputed and collectively stored in a “template bank”. The parameters of the templates are $\vec{\theta}=\{m1,m2,\chi_{1},\chi_{2}\}$ and in the detector frame Krolak and Schutz (1987).
To identify a real gravitational-wave signal $h$, we compute the correlation between the template waveform $h_{T,\;j}$ and the data $d$ (which are composed of the background noise and potentially a signal) to construct the SNR $\rho$. In the time domain, following Allen et al. (2012), this is
@@
The SNR $\rho(t)$ is defined as the modulus of the complex SNR, i.e., $\rho(t)=|z_{j}(t)|$.
@@
Figure 1: Left: match between lensed signals and unlensed templates $\mathcal{M}(h_{\rm{L}},h_{\rm{UL}})$ as a function of the redshifted lens mass $M_{\rm Lz}$ and impact parameter $y$. Right: ratio of the optimal signal-to-noise ratio of lensed $\rho_{\rm{L,opt}}$) and unlensed ($\rho_{\rm{UL,opt}}$) signals for the same $(M_{\rm Lz},y)$ parameter space). Here, we use a spinless binary black hole signal at 125 Mpc with $m_{1}=m_{2}=30M_{\odot}$. $\rho_{\rm{L,opt}}/\rho_{\rm{UL,opt}}>1$ holds true for point mass lens model.
@@ II.2 Current efforts in detectability study
In addition, the phase and time shifts, $\phi_{0}$ and $\tau_{0}$, maximize the match between the template waveform $h_{T,\;j}$ and the signal $h$. High $\mathcal{M}$ values indicate that $h_{T,\;j}$ and $h$ are well matched, where a maximum value of $\mathcal{M}=1$ indicates that the signal is exactly described by the template waveform. In the case where the template waveform exactly matches the signal (i.e. $h=h_{T,\;j}$) and the SNR is optimized with respect to noise realizations, the SNR is called the “optimal SNR” $\rho_{\rm opt}$ , and the “observed SNR” $\rho_{\rm obs}$ otherwise.
Note that $\mathcal{M}$ is normalized by the waveform amplitude in Eq. 5, whereas the SNR, as defined in Eq. 3, depends on both the waveform amplitude and waveform distortion. Consequently, $\mathcal{M}$ is sensitive to additional physics that alter the waveform morphology but is insensitive to amplitude changes. On the other hand, the SNR is sensitive to both waveform distortion and changes in amplitude.
@@ 6 Unlensed 2,3 1250 6217
Table 1: Set of 11 injections used in our three simulation campaigns. Injection campaign 1 studies the impact of lensing on detectability. Injection campaign 2 studies the role of signal strength. Injection campaign 3 gives a general picture of the detectability of the lensed gravitational waves. The found injection is defined as the FAR of the injection below the FAR threshold $3.85\times 10^{-7}$ Hz (see Sec. III.2 for FAR definition). All injections used the $30-30M_{\odot}$ spinless binary black hole waveform (see Sec. V.2 for details on other source parameters).
Various methods, each with its own strengths and limitations, have been used to assess gravitational-wave signal detectability. One widely used approach, employed in studies by Tiwari (2018); Veske et al. (2021); Gerosa et al. (2020), calculates the optimal SNR from a network of detectors, assuming that the same source parameters are recovered in matched-filtering searches. This method has been applied to evaluate the detectability of gravitational-wave signals with wave optics effects, as seen in works by Bondarescu et al. (2023); Mishra et al. (2024). While this assumption is not strictly valid due to noise fluctuations, it is a useful proxy for assessing detectability, particularly in high-SNR scenarios. Reference  Essick (2023) improved this method by developing semianalytic model to estimate the sensitivity of gravitational-wave searches by analytically deriving the distribution of matched-filter SNR in Gaussian noise, incorporating effects like template bank correlations and network SNR maximization. This approach enables rapid predictions of detection probabilities without requiring large-scale injection campaigns, and offers a computationally efficient solution for the study of detection rates of CBC sources that require analyzing millions of signals.
In matched-filtering search pipelines, computation of SNR is followed by methods to mitigate the fluence of non-Gaussian noise flucuation and compute the ranking statistics. For example, a signal consistency test has been employed to ensure the observed signals are consistent with the template bank. If the gravitational-wave signal contains addition physics that deviate from the template, such as eccentricity, higher mode, non-GR effects, gravitational lensing, etc, it is regarded as inconsistent with the template and results in a lower detection efficiency. Detailed discussion of the signal-consistency test with example is presented in Sec. III.
Since matched-filtering search pipeline is critical in GW detections, an inacccurate modelling of the detection lead to selection biases of population constraints. For examples,  Magee et al. (2024) attempted to consider the effects of signal consistency test to investigate detectability of beyond-GR effects and the selection biases induced and how they influence the population constraints. For the most comprehensive analysis of detectability, complete injection campaigns are necessary, which helps to investigate the impact of eccentricity Ramos-Buades et al. (2020) and higher-order modes Capano et al. (2014) on matched-filtering searches. These campaigns simulate the detection process using search pipelines and simulated gravitational-wave signals. This approach provides detailed information about the injected signals, including matched-filter SNR, signal-consistency test values, and final detection statistics. While computationally intensive, injection campaigns offer the most accurate and complete assessment of gravitational-wave signal detectability across various scenarios and detector configurations Messick et al. (2017).